### v1

# In-Model Merging for Enhancing the Robustness of Medical Imaging Classification Models

Abstract
Model merging is an effective strategy to merge multiple models for enhancing model performances, and more efficient than ensemble learning as it will not introduce extra computation into inference. However, limited research explores if the merging process can occur within one model and enhance the model’s robustness, which is particularly critical in the medical image domain. In the paper, we are the first to propose in-model merging (InMerge), a novel approach that enhances the model’s robustness by selectively merging similar convolutional kernels in the deep layers of a single convolutional neural network (CNN) during the training process for classification. We also analytically reveal important characteristics that affect how in-model merging should be performed, serving as an insightful reference for the community. We demonstrate the feasibility and effectiveness of this technique for different CNN architectures on 4 prevalent datasets. The proposed InMerge-trained model surpasses the typically-trained model by a substantial margin. The code will be made public.
Keywords:
Model Merging In-Model Merging Medical Imaging.
1 Introduction
Deep learning has become the cornerstone of modern machine learning. Ensemble learning is a effective manner for enhancing the model performances of medical imaging analysis (MIA), where tasks such as disease diagnosis demand high levels of accuracy. However, the inference computation will increase linearly.
Model merging, on the other hand, a strategy typically applied across multiple neural networks, has recently gained attention for its ability to improve model performance without introducing extra computation in the inference. By averaging weights from different models, Model Soups [20, 18] can lead to improved generalization and robustness with simply merging models. Ties-merging [21] takes one step ahead by taking weight conflicts into account while merging. There are multiple models perform merging from a different perspective by considering the merging process of finding a good way to interpolate on the loss basins [1, 8, 12, 15]. Existing works [20, 2, 21, 7, 3, 15] have demonstrated the potential of merging techniques to optimize neural network performance across multiple models. However, these techniques are inter-model, requiring multiple models trained for merging. An important research question remains and limited research has explored it: can we conduct merging operations within a single model to enhance model robustness but without introducing extra inference computation cost?
To answer this question and bridge the gap, in this work, we propose In-model Merging, a novel technique that selectively merges similar convolutional kernels within a single model. To the best of our knowledge, we are the first to consider merging techniques for a single model. By strategically merging similar kernels, the patterns redundancy between kernels can be diminished, a regularization effect will thus contribute to enhance the model’s robustness. We analytically demonstrate that shallow layers are crucial for low-level feature extraction and inappropriate for in-model merging. Also, we identify suitable similarity thresholds and merge weights/probabilities with extensive analysis experiments. Our work is generic, but it has important applications in MIA, which is our focus in the paper and where model robustness is essential. Our contributions are summarized as follows:
  • We propose a novel In-model Merging (InMerge) technique for enhancing the robustness of medical imaging classification models. To the best of our knowledge, we are the first to consider merging techniques from and within a single model.
  • We demonstrate how the proposed In-model Merging model can be integrated into different CNN networks as a plug-and-play module. We also show how this method enhances the model performance with minimal training cost and without introducing additional inference costs.
  • We demonstrate the effectiveness of the proposed In-model Merging in the challenging context of MIA. Through extensive analysis, this paper also reveals important characteristics that affect how In-model Merging should be performed to get useful insights.
2 Methodology
To improve the robustness of feature representation in CNNs, we propose a novel In-model Merging strategy. The model selectively merges similar convolutional kernels in the deeper layers of the model, while preserving the integrity of shallow layers. By merging kernels and reducing the kernel redundancies, the approach enhances model performance. It is worth noting that the proposed in-model merging method works as a finetuning approach. After a model is pretrained, In-model Merging requires a few epochs of finetuning, which ensures it will not introduce much computation overhead during the training process. In addition, no extra computation is added at inference time, as there will not be any merging operations at inference.
2.1 Kernel Similarity Computation
Figure 1: An example showing the similarity differences between patterns stored in kernels in a well-trained ResNet34 layer. The more similar pair of kernels would incur larger similarity. Kernel 9 and Kernel 10 are similar, the similarity $\text{sim}(\mathbf{k}_{9},\mathbf{k}_{10})=0.9372$; Kernel 3 and Kernel 11 nearly have orthogonal textures, so $\text{sim}(\mathbf{k}_{3},\mathbf{k}_{11})=-0.0234$; Kernel 4 and Kernel 12 have very different textures and colors, $\text{sim}(\mathbf{k}_{4},\mathbf{k}_{12})=-0.5049$; Kernel 5 and Kernel 13 are somewhat similar, $\text{sim}(\mathbf{k}_{5},\mathbf{k}_{13})=0.4562.$
The similarity between two convolutional kernels is measured using cosine similarity, which quantifies the angular difference in their vectorized representations. Given two kernels $\mathbf{K}_{i}$ and $\mathbf{K}_{j}$ from the same convolutional layer, we reshape them into one-dimensional vectors:
$\small\mathbf{k}_{i}=\text{vec}(\mathbf{K}_{i}),\quad\mathbf{k}_{j}=\text{vec}(\mathbf{K}_{j}),$ (1)
where $\text{vec}(\cdot)$ denotes the vectorization operation that flattens the kernel into a column vector. The cosine similarity is then computed as:
$\small\text{sim}(\mathbf{k}_{i},\mathbf{k}_{j})=\frac{\mathbf{k}_{i}^{T}\mathbf{k}_{j}}{{\|\mathbf{k}_{i}\|\|\mathbf{k}_{j}\|}},$ (2)
where |||||$\cdot$||| | ⋅ | | represents the $\ell_{2}$-norm. This metric ensures that similarity is independent of the absolute magnitudes of the kernels, focusing solely on their directional alignment.
2.2 In-Model Kernel Merging Strategy
To preserve the integrity of shallow layer features, the merging process is constrained to be only performed on deeper layers. In those convolutional layers, beyond the first $L_{s}$ shallow layers, similar kernels are merged dynamically during training. Let $\mathbf{W}={\mathbf{K}_{1},\mathbf{K}_{2},\dots,\mathbf{K}_{n}}$ denote the set of kernels in a convolutional layer. For each kernel $\mathbf{K}_{i}$, another kernel $\mathbf{K}_{j}$ is randomly selected such that $j\neq i$. If the similarity criterion is met, the two kernels are merged using a weighted interpolation with a certain probability $p$:
$\small\mathbf{K}_{i}\leftarrow\alpha\mathbf{K}_{i}+(1-\alpha)\mathbf{K}_{j},\;\;s.t.\;\text{sim}(\mathbf{k}_{i},\mathbf{k}_{j})>\tau,$ (3)
$\small P(\mathbf{K}_{i}\leftarrow\alpha\mathbf{K}_{i}+(1-\alpha)\mathbf{K}_{j})=p,$ (4)
where $\tau\in[-1,1]$ is the similarity threshold; $\alpha\in[0,1]$ is the weighting factor that controls the balance between the two kernels. This merging operation helps mitigate redundancy in feature extraction while preserving the network expressiveness. This probabilistic merging mechanism prevents the network from over-collapsing to a limited set of feature extractors, giving the effect of regularization, thereby retaining sufficient expressiveness while enhancing generalization.
2.3 Algorithm and Rationale
The In-model Merging algorithm is shown as follows:
Input: Loaded pretrained model $\mathcal{M}$ with convolutional layers $\{\mathcal{L}_{1},\mathcal{L}_{2},\dots,\mathcal{L}_{N}\}$, shallow layer count $L_{s}$, merging probability $p$, similarity threshold $\tau$.
Output: Updated model $\mathcal{M}$ with in-model merging.
for each training iteration do
       for each convolutional layer $\mathcal{L}_{i}$ in $\mathcal{M}$ do
             if $i<L_{s}$ then
                   continue;
            Retrieve kernel weights $\mathbf{W}=\{\mathbf{K}_{1},\mathbf{K}_{2},\dots,\mathbf{K}_{n}\}$;
             Initialize new weights $\mathbf{W}^{\prime}\leftarrow\mathbf{W}$;
            for each kernel $\mathbf{K}_{i}\in\mathbf{W}$ do
                   With probability $p$, randomly select another kernel $\mathbf{K}_{j}$ such that $j\neq i$;
                  Compute cosine similarity:
$\text{sim}(\mathbf{k}_{i},\mathbf{k}_{j})=\frac{\mathbf{k}_{i}\cdot\mathbf{k}_{j}}{\|\mathbf{k}_{i}\|\|\mathbf{k}_{j}\|}$
                  if $\text{sim}(\mathbf{k}_{i},\mathbf{k}_{j})>\tau$ then
                         Merge kernels:
$\mathbf{K}_{i}\leftarrow\alpha\mathbf{K}_{i}+(1-\alpha)\mathbf{K}_{j}$
Update $\mathbf{W}^{\prime}[i]\leftarrow\mathbf{K}_{i}$;
            Update layer weights via back-propagation: $\mathcal{L}_{i}.\mathbf{W}\leftarrow\mathbf{W}^{\prime}$;
Algorithm 1 In-model Merging for Enhancing the Robustness of Medical Imaging Classification Models
As an example shown in Fig. 1, the similar kernel pair has a larger similarity, while the dis-similar kernel pair has a smaller similarity. The rationale is that there are patterns stored in learned weights. Convolutional kernels in different layers often capture overlapping or similar patterns, which introduce redundancy in learned features. When merging the similar patterns, it reduces redundant components within one single model. The process therefore has the same effect of regularization, so that the redundancy can be suppressed and model robustness can be enhanced. This selective and iterative merging approach ensures the model maintains a balance between feature diversity and redundancy reduction.
3 Experiments
3.1 Datasets and Implementation Details
ChestXRay14 Tasks ChestXRay-14 is a chest X-ray imaging dataset for multi-label classification. It contains 112,120 frontal chest X-ray images from 30,805 unique patients. The dataset is labeled with 14 different lung disease categories. For the dataset, we adopt the official split for training and evaluation. For model training, we apply a series of standard data augmentations and transformations. These include random horizontal flipping, resizing the images to 224 by 224 pixels. The model is trained with a sigmoid head on 14 outputs and using 14 Binary Cross-Entropy Losses for the multi-label classification task. The optimizer chosen for the training process is Stochastic Gradient Descent (SGD) with a staged learning rate strategy and an initial learning rate of 0.01, momentum set to 0.9, and weight decay of 1e-4 for regularization. The batch size is set to 324. The default backbone network is based on ImageNet pretrained VGG19 model. After the backbone model is trained (by default 30 epochs), an extra 10 epochs for in-model merging finetuning is applied.
MedMNIST Tasks MedMNIST data is more diverse to the chestxray data. PathMNIST, DermaMNIST, and OCTMNIST are in the collection of MedMNIST and commonly used in MIA. PathMNIST is a dataset of histopathological images with 9 classes, particularly focused on classifying tissue samples, such as breast cancer tissue, into categories like malignant or benign. DermaMNIST, on the other hand, contains dermoscopic images of skin lesions with 7 classes and is typically used for the classification of various skin diseases, such as melanoma and non-melanoma types. OCTMNIST consists of optical coherence tomography (OCT) images of the retina with 8 classes and is mainly used for detecting different retinal diseases. The implementation fine-tunes a pre-trained VGG19 model on the MedMNIST dataset. Data preprocessing includes resizing images to 64 by 64 and using a batch size of 1024. Training (20 epochs backbone training and 5 epochs InMerge training) is performed using SGD with a multi-step learning rate schedule. The model’s performance is evaluated on a validation set each epoch, with the best-performing model saved. In the experiments, each model is trained and tested 5 times to fetch its mean and standard deviation.
For all these datasets, we generally set merging weight $\alpha$ to 0.8; skipped layer $L_{s}$ to 10 for VGG19 as default; merging probability $p$ to 0.3 and similarity threshold $\tau$ to 0.3. In-model Merging is abbreviated as InMerge in the following figures/tables/descriptions. To maintain a fair comparison, the baseline models on all datasets are trained with the same number of epochs as the InMerge model, and we focus our analysis on CNN models.
3.2 Results & Discussion
Table 1: Performance Comparison of Different Models in AUROC. The 14 different lung disease categories are Cardiomegaly, Effusion, Infiltration, Mass, Nodule, Pneumonia, Pneumothorax, Consolidation, Edema, Emphysema, Fibrosis, Pleural Thickening, Hernia (the abbreviations in the table keep this order). “DualCXN” in the table denotes DualCheXNet. The best-performed model for each column is bolded.
Models Atel Card Effu Infi Mass Nodu Pneu1 Pneu2 Cons Edem Emph Fibr P_T Hern Ave
U-DCNN [19] 0.700 0.810 0.759 0.661 0.693 0.669 0.658 0.799 0.703 0.805 0.833 0.786 0.684 0.872 0.745
LSTM [22] 0.733 0.856 0.806 0.673 0.718 0.777 0.684 0.805 0.711 0.806 0.842 0.743 0.724 0.775 0.761
AGCL [16] 0.756 0.887 0.819 0.689 0.814 0.755 0.729 0.850 0.728 0.848 0.906 0.818 0.765 0.875 0.803
CheXNet [14] 0.769 0.885 0.825 0.694 0.824 0.759 0.715 0.852 0.745 0.842 0.906 0.821 0.766 0.901 0.807
DNet [11] 0.767 0.883 0.828 0.709 0.821 0.758 0.731 0.846 0.745 0.835 0.895 0.818 0.761 0.896 0.807
CRAL [10] 0.781 0.880 0.829 0.702 0.834 0.773 0.729 0.857 0.754 0.850 0.908 0.830 0.778 0.917 0.816
CAN [13] 0.777 0.894 0.829 0.696 0.838 0.771 0.722 0.862 0.750 0.846 0.908 0.827 0.779 0.934 0.817
DualCXN [4] 0.784 0.888 0.831 0.705 0.838 0.796 0.727 0.876 0.746 0.852 0.942 0.837 0.796 0.912 0.823
LLAGnet [6] 0.783 0.885 0.834 0.703 0.841 0.790 0.729 0.877 0.754 0.851 0.939 0.832 0.798 0.916 0.824
CheXGCN [5] 0.786 0.893 0.832 0.699 0.840 0.800 0.739 0.876 0.751 0.850 0.944 0.834 0.795 0.929 0.826
InMerge (ours) 0.805 0.895 0.881 0.710 0.842 0.750 0.770 0.878 0.801 0.888 0.915 0.818 0.773 0.903 0.830
Chest X-Ray Tasks As shown in Tab. 1, the proposed InMerge model consistently outperforms existing techniques across multiple disease categories, achieving the highest average performance. Notably, it ranks first in 10 out of 15 ranks (including average), demonstrating its broad applicability and robustness. Particularly, InMerge exhibits large improvements in conditions such as Atelectasis (0.805), Effusion (0.881), and Consolidation (0.801) in AUC, suggesting that its ability to refine feature representation through in-model merging is highly beneficial for identifying complex patterns of these diseases. These results indicate that selectively merging similar convolutional kernels in deeper layers enhances representation learning with minimal computational burden.
Compared to prior methods, InMerge provides a noticeable improvement over current best architectures, e.g., CheXGCN, which shows strong performance in most cases but falls short in several other conditions. While CheXGCN benefits from graph-based relational modeling, InMerge consistently maintains high accuracy across diverse disease categories. Additionally, models such as CAN, DualCheXNet and LLAGnet, which integrate multi-branch or attention-based mechanisms, demonstrate competitive performance but do not consistently surpass InMerge, indicating that incorporating more parameters in the training may not always be the most effective strategy. Furthermore, when examining disease categories such as Pneumonia and Edema, InMerge also exhibits strong results. The strength of InMerge lies in its ability to dynamically refine internal representations by merging redundant yet functionally similar convolutional kernels, leading to improved generalization without the need for explicit architectural modifications or additional supervision.
Table 2: Performance comparison of Baseline and InMerge across different MedMNIST datasets in accuracy.
Dataset Cls# Samp# Train/Val/Test Baseline InMerge (ours)
PathMNIST 9 107,180 89,996/10,004/7,180 0.898±0.010 0.924±0.002
DermaMNIST 7 10,015 7,007/1,003/2,005 0.745±0.009 0.760±0.005
BloodMNIST 8 17,092 11,959/1,712/3,421 0.941±0.005 0.948±0.004
MedMNIST Tasks In Tab. 2, we present a comparative classification accuracy evaluation of our proposed In-Model Merging (InMerge) method against the baseline (trained with the same total number of epochs with the InMerge model) across multiple MedMNIST datasets, covering diverse medical imaging tasks. The results consistently demonstrate that InMerge outperforms the baseline across all datasets. These improvements are particularly noteworthy given that InMerge introduces minimal architectural modifications and computational costs. In terms of standard deviation, the results of In-model Merging across the three datasets consistently exhibit lower values compared to its baseline model, demonstrating greater stability across different training runs.
The best improvement is observed in PathMNIST, where InMerge achieves an average accuracy of 92.4%, surpassing the baseline’s 89.8% by a large margin (2.9% improvement). It suggests that our method effectively enhances feature representation in complex tissue patterns via regularization in convolutional kernels, leading to more robust feature extraction. Similarly, in BloodMNIST, involving microscopic blood cell images, InMerge yields a notable improvement. The enhancement on DermaMNIST (76.0% vs. 74.5%) is meaningful in dermatology applications where subtle textural differences are critical for classification. This suggests that even in datasets with limited training samples, InMerge merges redundant convolutional kernels to provide a regularization effect, reducing overfitting while preserving critical feature diversity. These results suggest the robustness of In-model Merging for medical image classification tasks.
3.3 Analyses
The analysis experiments are conducted on the chest X-ray data.
Merging layer $L_{s}$ sensitivity Fig. 2(a) shows performance variations across merging different layers ($\geq$ Layer 10 means any layer below 10th layer is not considered for merging). A clear trend shows where only including deeper layers (Layer 10 and Layer 15) outperform including from shallower ones (Layer 0 and Layer 5) in most categories. This highlights the importance of deep feature representations for disease identification. Notably, conditions such as Emphysema, Cardio., and Pneumo. show substantial improvements when merging deeper layers, suggesting the possibility that these conditions rely more on low-level abstract features, so they are sensitive to layer change. For the setup of layer 0 and 5, the model performances are very different from layer 10, but layer 10 and 15 are not very different, showing the general stability when layers go deeper. These findings empirically prove the idea that deeper feature merging is crucial for model merging strategies, while also indicating that the optimal layer depth for different diseases may vary, which is valuable for task-specific model merging setup.
Merging with different similarities $\tau$. Fig. 2(b) compares model performance under varying similarity thresholds for in-model kernel merging. A clear trend emerges where higher similarity thresholds (e.g., $\geq$ Sim 0.3) generally lead to stronger performance than the thresholds 0.1 or Dissimilar condition, where more dissimilar kernels are merged, indicating that merging dissimilar kernels disrupts critical feature representations. Interestingly, for Effusion and Edema, the performance difference between different similarity settings is relatively small, suggesting that these conditions may be less sensitive to in-model merging. These results highlight the importance of selective merging strategies, where preserving high-similarity features is beneficial while merging dissimilar kernels can degrade performance.
Ablation Study, and Merge with different merge weights $\alpha$ / probabilities $p$. From Fig. 2(c), we can observe from the perspective of either different merge weights $\alpha$ or probabilities $p$, the line plots show a stable trend and highly above the baseline model (trained with the same number of epochs of InMerge), also serving as the ablation study of w and w/o InMerge.
In-model Merging on different CNN architectures. We also examine the generality of In-model Merging on different models. For VGG16 model ($L_{s}=$10), the average AUC of In-model Merging surpasses the baseline model (0.817 VS. 0.807). Similar results are shown on ResNet34 ($L_{s}=$30), the model performance gained from 0.792 to 0.796.
(a) Layer thresholds.
(b) Different similarities.
(c) Merge weight + prob.
Figure 2: The sensitivity of different merged layers (a), merge with different similarities (b), merge with different merge weights and merge probabilities (c).
4 Conclusion
We propose In-model Merging, where the merging process occurs within one model and enhances the model robustness with minimal training computations and no additional computations in inference. By selectively merging similar kernels in deeper layers, the method enhances feature representation while preserving low-level features for robustness. Our experiments confirm the effectiveness of in-model merging in improving model performance and robustness. Results across multiple medical datasets show improved generalization. This suggests that in-model merging is a practical enhancement for neural networks with potential for further exploration. Future research may focus on exploring InMerge in different network architectures, such as Transformers [17] or Mamba [9].